OpenStudy (anonymous):

JANE, ANDRE, AND MARIA PICK APLLES. ANDRE PICKS THREE TIMES AS MANY POUNDS AS MARIA. JANE PICKS TWO TIMES AS MANY POUNDS AS ANDRE. THE TOTAL WEIGHT OF THE APPLES IS 840. HOW MANY POUNDS OF APPLES DOES ANDRE PICK?

OpenStudy (anonymous):

Okay so let x = the amount of apples that Maria picks. Andre picks three times as much as Maria, so amount Andre picks = 3x Jane picks two times as much as Andre, and Andre = 3x So Jane = 2(3x), = 6x All of these added together should give 840. x + 3x + 6x = 840 10x = 840. x = 84 Therefore, Maria picked 84 pounds of apples Andre picked 3x = 3 x 84 = 252 pounds of apples. Jane picked 6x = 6 x 84 = 504 pounds of apples.

OpenStudy (anonymous):

Just to make sure the answer is correct, you can add all the pounds together. 84 + 252 + 504 = 840 So the answer is correct.